Calculate the number of sulfur atoms of 150 g in the antibiotic cefdinir
1
Expert's answer
2019-10-15T08:36:04-0400

Molecular formular of cefdinir is "C_{14}H_{13}N_5O_5S_2"

Mollar mass = "395.4 \frac{g}{mol}"

"n(cefdinir)=\frac{m}{M} = \frac{150 g}{395.4\frac{g}{mol}}=0.379 mol"

As one molecule of sefdinir contains 2 atoms of S then n(S)= 2*0.379 mol = 0.758 mol


"N(S)=n(S)\times N_A= 0.758\times 6.022\times 10^{23} = 4.57\times 10^{23 }atoms"